OpenStudy (anonymous):

Solve the absolute value inequality: |4x + 3| 19

Parth (parthkohli):

\(> \text{ or} <\)?

Parth (parthkohli):

Or possibly \(\le\) or \(\ge\)?

OpenStudy (anonymous):

>=

Parth (parthkohli):

Aha. Then \(4x + 3\)'s distance [geometrically] is 19 or more from 0.

Parth (parthkohli):

Therefore, \(4x + 3 \ge 19\) or \(4x + 3 \le -19\)

Parth (parthkohli):

do you get it? Now you may solve for \(x\) in both inequalities.

OpenStudy (anonymous):

yeah go on

Parth (parthkohli):

What else am I missing in my explanation? I believe that it is already completed.

OpenStudy (anonymous):

oh ok wel then if i go follown ure way my answer becomes x <= -5.5 and x >= 4

Parth (parthkohli):

You are correct :)

OpenStudy (anonymous):

thanks for ure help :0

Parth (parthkohli):

yw